# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v3

@@ 4.1 Experimental Setting
Evaluation Metrics. We adopt the same evaluation metrics used in prior works, which include “Transfer”, “Average”, and “Last”. The “Transfer” metric evaluates the model’s zero-shot transfer capability on previously unseen data. The “Last” metric assesses the model’s ability to retain knowledge from earlier tasks. The “Average” metric is a composite score that reflects the mean performance across both the “Transfer” and “Last” evaluations.
Implementation Details. We follow the experimental setups in [62, 57, 54] and utilize the CLIP model with a ViT-B/16 backbone [39] for all experiments. By default, our proposed importance-aware sparse LoRA is applied to all pre-trained weight matrices across both the vision encoder and text encoder, with an initial rank set to 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [29]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold value for adaptive pruning is fixed at 0.005 for both settings.
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [39] 68.24 33.77 66.72 56.24
MoE-Adapter [57] 68.24 33.77 66.72 56.24
RAIL [54] 68.24 33.77 66.72 56.24
Ours 68.95 36.52 68.52 58.00
Table 3: Evaluation of average transfer accuracy on datasets that were never encountered during the continual learning stages.
@@ Methods
Ours 4.4M None
Table 4: Comparison on computation cost with previous methods.
@@ 4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the evaluation results on the X-TAIL setting in Table 1. Unlike previous works that rely on model parameter updates [25, 51, 40, 13] along with additional memory replay or reference data, our approach performs parameter-efficient updates without requiring any memory replay or reference data. This results in significant improvements across all evaluated metrics.
In contrast to prior methods [57, 54] [^2] Dual version of [54] stores all the training samples’ feature embeddings in the memory, for a fair comparison, we excluded the results here. We leave more detailed comparisons in Appendix., which use zero-shot predictions to preserve pre-trained capabilities and prevent forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are constrained by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly impacted due to the nature of continual parameter updates, our method ultimately achieves a higher overall transfer accuracy by effectively adapting the pre-trained model to new tasks while also enhancing its generalization capabilities.
For both the Average and Last metrics, our method demonstrates superior or comparable performance without relying on domain ID prediction [57], high-dimensional projections, or storing a memory bank of previously seen sample features [54]. Instead, our method focuses purely on updating the pre-trained model, yet still delivers competitive results.
Furthermore, since our approach is entirely orthogonal to the additional techniques used in previous methods [57, 54], these techniques can be seamlessly integrated into our framework to further enhance performance, benefited from our ultimate goal of continually enhancing the pre-trained model’s capabilities.
Multi-domain task-incremental learning. We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [57, 54]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this setting. Without employing any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average Transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works. Moreover, by supplementing our method with techniques used in prior approaches [57, 54], we can further improve the performance.
@@ 4.3 Discussions
Pre-trained knowledge retention and potentially improved pre-trained model. To understand how our method improves transfer performance beyond the upperbound set by zero-shot predictions in previous works, we evaluate on datasets that were never encountered during the continual learning steps, specifically CIFAR100 [24], Places365 [63], and ImageNet-1k [10], as shown in Table 3.
@@
Unlike prior approaches [57, 54] that rely exclusively on zero-shot predictions for unseen data, thereby limiting their performance to that of the original CLIP model, our method goes beyond this limitation. By effectively integrating new knowledge into the pre-trained parameter space without sacrificing existing capabilities, our approach enhances the model’s overall performance. Consequently, our method achieves superior transfer accuracy, surpassing the theoretical upper bound defined by zero-shot performance in previous methods. This not only demonstrates improved adaptation but also enhances the model’s generalization capabilities on completely unseen datasets during CL.
@@
Rank allocation results on different datasets. We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig. 6 and Fig. 6. Overall, it is evident that the Aircraft dataset generally requires more ranks to be updated in the pre-trained model compared to the Pets dataset.
@@
We observe that for the Aircraft dataset, a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig. 6(b)). Additionally, a higher concentration of ranks is assigned to the deeper layers, particularly layers 6 to 10 (Fig. 6(c)).
@@
In contrast, the analysis of the Pets dataset (Fig. 6(b) and Fig. 6(c)) reveals that more ranks are allocated to the Attentions’ Key and Value projection, and MLP’s FC layer of vision encoder. Furthermore, the majority of rank allocations are in the first and last layers, suggesting a different pattern for promoting knowledge retention and adaptation across datasets. These findings highlight the adaptive nature of our approach, where rank allocation varies based on the characteristics of the dataset, optimizing the model’s ability to capture domain-specific information effectively.
@@
Computation Cost. We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig. 4. For MoE-Adapter, the difference between training and testing parameters stems from its mixture-of-experts design: during training, all experts are active, while at inference, only the top-2 experts are selected, reducing the active parameter count. RAIL-Primal incurs additional computational overhead during inference due to its use of high-dimensional feature expansion through projection and regression layers. RAIL-Dual maintains a memory bank of image features for all training samples across 1,100 classes, which significantly increases memory usage.
Notably, these methods experience linearly growing computational costs as the number of tasks increases. In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance. Once training is complete, these updates are merged into the pre-trained weights, resulting in zero additional memory overhead during inference.
@@ 4.4 Ablation Studies
Effects of insertion locations and number of initial ranks. We examine the impact of different insertion locations and initial ranks for our method, with results presented in Fig. 7. These findings are consistent with our analytical experiments shown in Fig. 3. Our results indicate that when training is restricted to only the vision encoder, transfer accuracy is significantly reduced, suggesting that the pre-trained capabilities are negatively impacted. Conversely, increasing the maximum rank improves adaptation ability without necessarily reducing transfer accuracy. Furthermore, when limiting insertion to only the Attention modules, we observe that transfer accuracy can surpass our previously reported state-of-the-art results in Table 1 and Table 2. However, this comes at the expense of reduced adaptation capabilities. To achieve an better balance between adaptation and retention, we enable insertion into all modules by default. This configuration ensures a better trade-off between adapting to new tasks while preserving the model’s zero-shot transfer capabilities.
Effect of maximum threshold value and ratio for dense iterations. We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio for dense training iterations in Fig. 8. The results indicate that as the threshold increases, the Transfer accuracy shows an increasing trend. This suggests that pruning more ranks can enhance the model’s performance on unseen data (i.e., improve transfer capability), but it also limits the model’s ability to adapt to downstream tasks, as evidenced by changes in the Last accuracy metric. Our experiments show that a threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
The ratio of dense training iterations determines when the model begins the adaptive pruning process. Setting this ratio to 0 (i.e., starting pruning from the beginning of training) severely hinders adaptation, as LoRA lacks sufficient time to learn essential task-related knowledge before pruning begins. On the other hand, delaying pruning for too long can result in the model forgetting pre-trained capabilities, thereby reducing transfer accuracy. Our findings indicate that setting the ratio around 0.5 to 0.7 works best across all experimental settings, providing a good balance between adaptation and knowledge retention.
@@ 5 Conclusion
In this paper, we introduced a novel Dynamic Rank-Selective LoRA approach to address the challenges of continual learning for vision-language models. By dynamically adjusting the ranks based on each ranks’ importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead. Extensive evaluations on multi-domain and cross-domain settings demonstrated that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and zero-shot generalization.
Limitations and future works. Our current approach directly merges parameter updates into the pre-trained weights without explicitly considering correlations between tasks. Exploring methods to leverage inter-task relationships and shared information build a promising direction for future research.
@@ B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
Related continual learning methods focus on updating representations through the backbone model (e.g., [50, 62, 57], and ours) or by adding a representation alignment or adaptive prediction head [33, 54]. These two approaches can be considered orthogonal. We investigate how the representation updates introduced by our proposed method can complement the additional adaptive prediction head, utilizing the aggression-based adapters described in [54].
@@
Since RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model. In contrast, our approach gradually accumulates knowledge into the pre-trained model, enhancing its capability on unseen data and achieving improved Transfer accuracy. Moreover, by leveraging our continually enhanced pre-trained model, coupling our method with RAIL significantly outperforms the original RAIL, which uses a frozen pre-trained model, on the Average and Last metrics.
@@
Ours + RAIL-Primal 63.1 74.1 83.4
@@
Ours + RAIL-Dual 63.1 74.2 83.6
@@
$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{T}\mathbf{W}\mathbf{V}^{T}\|_{F}},$ (8)
where $\mathbf{W}$ is the pre-trained weight matrix, and $\Delta\mathbf{W}$ is the parameter change introduced by LoRA. For vanilla LoRA and our method, $\Delta\mathbf{W}$ is computed using Eq. (1) and Eq. (2), respectively. $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
@@
(d) Aircraft - Ours
(e) EuroSAT - Ours
(f) Pets - Ours
@@
B.3 More Visualizations and Analyses of Rank Allocation Across Datasets

### v4

@@ 4.1 Experimental Setting
Evaluation Metrics. We adopt the same evaluation metrics as prior works, including “Transfer”, “Average” , and “Last”. “Transfer”evaluates the model’s zero-shot transfer capability on unseen data. “Last” assesses the model’s ability to retain knowledge from earlier tasks, and“Average” provides a composite score, representing the mean performance across both the “Transfer” and “Last” evaluations.
Implementation Details. We follow the experimental setups in [63, 58, 55] and use the CLIP model with a ViT-B/16 backbone [40] for all experiments. By default, CoDyRA is applied to all pre-trained weight matrices in both the vision and text encoders, with an initial rank of 16. Each task in both experimental settings is trained over 500 iterations using the AdamW optimizer [31]. The dense training iteration ratio is set to 0.5 for X-TAIL and 0.7 for MTIL, relative to the total training iterations. The maximum threshold adaptive pruning is fixed at 0.005 for both settings.
@@ Methods
CoDyRA 4.4M None
Table 3: Comparison of computational cost with prior methods.
@@ 4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report the X-TAIL setting evaluation results in Table 1. In contrast to prior methods [58, 55] [^3] The DUAL version of [55] stores all the training samples’ feature embeddings in memory. For a fair comparison, we exclude these results here and provide more detailed comparisons in the Appendix., which rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, our method overcomes this limitation by continually refining the model. Notably, while these approaches are bounded by the upper limit of zero-shot performance, our approach surpasses this limit, achieving a higher average transfer accuracy. Although some predictions on unseen data might be slightly affected due to the continual parameter updates, our method ultimately improves overall transfer accuracy by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
[…]
Multi-domain task-incremental learning. We further evaluate our proposed method under the few-shot MTIL setting, as shown in Table 2, following [58, 55]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario. Without incorporating any additional techniques, our approach outperforms previous methods across all evaluation metrics, achieving an average transfer accuracy that surpasses the upper bound set by zero-shot predictions in prior works. Furthermore, by integrating techniques from previous approaches [58, 55], our method can be further enhanced, leading to even greater performance gains.
@@ 4.3 Discussions
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot [40] 68.24 33.77 66.72 56.24
MoE-Adapter [58] 68.24 33.77 66.72 56.24
RAIL [55] 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Table 4: Evaluation of average transfer accuracy on unseen datasets not encountered during continual learning.
@@
Pre-trained knowledge retention and potential pre-trained model improvement. To understand how our method improves transfer performance beyond the upper bound set by zero-shot predictions in prior works, we evaluated on hold-out datasets never encountered during continual learning, specifically CIFAR100 [25], Places365 [64], and ImageNet-1k [11], as shown in Table 4.
@@
Unlike prior approaches [58, 55], which rely exclusively on zero-shot predictions for unseen data—thereby constraining their performance to that of the original CLIP model—our method overcomes this limitation. By effectively integrating new knowledge into the pre-trained parameter space while preserving existing capabilities, our approach enhances overall model performance. As a result, it achieves superior zero-shot transfer accuracy, surpassing the zero-shot performance limits of prior works. This demonstrates that CoDyRA effectively incorporates general knowledge into its knowledge base, further improving its overall representation capability.
@@
Rank allocation across different datasets. We visualize and analyze the distribution of activated ranks merged into the pre-trained weights across different datasets in Fig. 6 and Fig. 6. Overall, the Aircraft dataset generally requires more rank updates in the PTM than the Pets dataset.
@@
For the Aircraft dataset, we observe that a greater number of ranks are allocated to the Output projection of the Attention module within the text encoder (Fig. 6(b)). Additionally, a higher concentration of ranks is assigned to deeper layers, particularly layers 6 to 10 (Fig. 6(c)).
@@
In contrast, for the Pets dataset, more ranks are allocated to the Attention module’s Value and Output projections and the MLP’s Projection layer within the vision encoder (Fig. 6(b)). Additionally, rank allocation is primarily concentrated in the first and last layers (Fig. 6(c)), indicating a distinct adaptation pattern that balances knowledge retention and domain adaptation across datasets. These observations highlight the adaptive nature of our approach, where rank allocation dynamically adjusts based on dataset characteristics, ensuring effective domain adaptation while preserving pre-trained knowledge. We provide more visualizations and analyses in Appendix B.3.
@@
Computation cost. We compare the computational costs in terms of trainable parameters and additional inference requirements, as summarized in Fig. 3. For MoE-Adapter, the difference between training and testing parameters arises from its mixture-of-experts design: during training, all experts are active, whereas during inference, only the top-2 experts are selected, reducing the number of active parameters. RAIL-Primal incurs additional inference overhead due to its use of high-dimensional feature expansion through projection and regression layers. Meanwhile, RAIL-DUAL significantly increases memory usage by maintaining a memory bank of image features for all training samples across 1,100 classes. Notably, these methods experience linearly growing computational costs as the number of tasks increases. In contrast, our method minimizes training overhead by selectively retaining only ranks with high importance. Once training is complete, these updates are merged into the pre-trained weights, ensuring zero additional memory overhead during inference.
@@ 4.4 Ablation Studies
Effects of insertion locations and number of initial ranks. We examine the impact of different insertion locations and initial ranks in our method, with results presented in Fig. 7. These findings are consistent with our analyses shown in Fig. 4. Our results indicate that when training is restricted to the vision encoder, transfer accuracy drops significantly, suggesting that the pre-trained capabilities are negatively impacted. Moreover, at specific locations, increasing the maximum rank improves adaptation ability while not always reducing transfer accuracy.
Interestingly, restricting updates to only the Attention modules can surpass our reported state-of-the-art Transfer metric in Table 1 and Table 2. However, this comes at the expense of reduced adaptation capabilities. To strike a better balance between adaptation and retention, we enable updates across all modules by default, ensuring an optimal trade-off between learning new tasks and retaining the model’s pre-trained capabilities.
Effect of maximum threshold value and dense iteration ratio. We analyze the effects of varying the maximum threshold value $\kappa_{\text{max}}$ and the ratio of dense training iterations in Fig. 8. Our results indicate that as the threshold increases, Last accuracy improves, suggesting that pruning more ranks can enhance transfer performance on unseen data. However, aggressive pruning limits adaption ability to to downstream tasks, as reflected in changes to the Last accuracy metric. We find that setting the threshold value around 0.005 strikes a good balance between retaining pre-trained knowledge and adapting to new tasks.
The dense training iteration ratio determines when adaptive pruning begins. Setting this ratio to 0 (i.e., pruning from the start of training) hinders adaptation, as the model lacks sufficient time to learn task-relevant knowledge before pruning begins. Conversely, delaying pruning too long increases the risk of forgetting pre-trained capabilities, ultimately reducing transfer accuracy. We find that setting this ratio between 0.5 and 0.7 achieves the best trade-off across all experimental settings, effectively balancing both adaptation and knowledge retention.
@@ 5 Conclusion
In this paper, we introduced CoDyRA, a novel approach to addressing continual learning challenges in vision-language models. By dynamically adjusting ranks based on their importance, our method optimizes knowledge retention while enabling efficient adaptation to new tasks without adding inference overhead. Extensive evaluations across various continual learning settings demonstrate that our method achieves state-of-the-art performance, effectively balancing downstream adaptation and pre-trained retention.
Limitations and future works. Our approach merges updates into pre-trained weights without modeling inter-task correlations. Future works could explore leveraging inter-task relationships and extending our method to other VLMs.
@@ B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
The related continual learning methods focus on updating the representation with the backbone model (e.g., [51, 63, 58], and CoDyRA) or newly added representation alignment or prediction head [35, 55]. These two types of methods can be seen as orthogonal. We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [55].
@@
RAIL is a regression-based method that applies an additional prediction head on top of the visual representations extracted by the frozen pre-trained vision encoder of the CLIP model. This approach is fully orthogonal to our method, as our focus is on continually incorporating new knowledge into the pre-trained model through dynamic low-rank parameter updates, and we can seamlessly integrate their techniques into our method. Because RAIL heavily relies on the frozen pre-trained model, the Transfer accuracy is limited by the zero-shot performance of the pre-trained model. In contrast, our approach gradually accumulates knowledge into the pre-trained model, enhancing its capability on unseen data and achieving improved Transfer accuracy. Moreover, by leveraging our continually enhanced pre-trained model, coupling our method with RAIL significantly outperforms the original RAIL, which uses a frozen pre-trained model, on the Average and Last metrics.
@@
CoDyRA + RAIL-Primal 63.1 74.1 83.4
@@
CoDyRA + RAIL-Dual 63.1 74.2 83.6
@@
$\text{Amp}=\frac{\|\Delta\mathbf{W}\|_{F}}{\|\mathbf{U}^{\text{T}}\mathbf{W}\mathbf{V}^{\text{T}}\|_{F}},$ (8)
where $\mathbf{W}$ is the pre-trained weight matrix, and $\Delta\mathbf{W}$ is the parameter change introduced by LoRA. For vanilla LoRA and CoDyRA, $\Delta\mathbf{W}$ is computed using Eq. (1) and Eq. (2), respectively. $\mathbf{U}$ and $\mathbf{V}$ represent the top $r$ singular vectors of $\text{SVD}(\Delta\mathbf{W})$, where $r$ is the number of interested ranks in analysis.
@@
(d) Aircraft - CoDyRA
(e) EuroSAT - CoDyRA
(f) Pets - CoDyRA
@@
B.3 More Visualizations and Analyses on Rank allocation results of different datasets
@@
B.4 Evaluation with Other VLMs
To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [26] on X-TAIL (Table 6). Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
Transfer Average Last
Zero-shot 51.03 46.45
[…]
Table 6: Averaged performance using BLIP.
B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
Leveraging our dynamic rank-adaptive parameter updates, our method naturally extends to Parameter-Efficient Fine-Tuning (PEFT) tasks. Table 7 presents 16-shot PEFT results, using the same hyperparameters as the main paper, showing that CoDyRA outperforms vanilla LoRA while reducing final trainable parameters by pruning ranks with zero importance. Notably, across multiple downstream tasks (Caltech101, Cars, and OxfordPet), CoDyRA achieves higher accuracy while using fewer parameters compared to LoRA with a fixed rank of 16.
These results highlight a more adaptive and efficient alternative to conventional low-rank tuning by automatically selecting the most relevant ranks based on training data. Crucially, by amplifying important ranks through learned higher importance weights, our approach significantly enhances downstream adaptation.
Acc. $\uparrow$ / Final Used Params. $\downarrow$
Caltech101 Cars OxfordPet
Zero-shot 88.40 64.70 89.00
LoRA ($r=16$) 96.43 / 4.4M 80.66 / 4.4M 92.94 / 4.4M
CoDyRA  ($r_{init}=16$) 97.20 / 3.95M 81.61 / 3.75M 94.44 / 3.85M
Table 7: Performance of 16-shot PEFT on downstream tasks.